A firms short run production function is given by Q=c * L + B * L^2 - a * L^3


at a=17.6, b=84.9 and c=73


determine the value of L which maximises the average product of labour rounding your answer to two (2) decimal places


1
Expert's answer
2021-08-20T08:57:23-0400

Let us define points where MP(L)="Q'(L)=0" .

MP(L)=c+2"\cdot B\cdot L-3\cdot a \cdot L^2=73+169.8\cdot L-52.8\cdot L^2=0" ;

This is a quadratic equation.

"D=169.8^2+4\cdot 73\cdot 52.8=44269.64"

"L_1={-169.8+\sqrt{44269.64}\over {-2\cdot 52.8}}=-0.384- extraneous\space root"

"L_2={-169.8-\sqrt{44269.64}\over {-2\cdot 52.8}}=3.60"

L2 is the maximum point of Q(L) because from graph of parabola Q'(L)

we see that Q'(L)>0 on (0, 3.6) and Q(L) increases;

and Q'(L)<0 on (3.6, "\infty)" Q(L) is decreasing.

So Lmax=3.60 is a value of factor L where average product is maximum.